Find each of the function below, indicate whether the function in onto, on-to-one neither or both. If the function is not onto or nor one-to-one, give an example showing why

                               

F;R      R. f(x)=x^2

                           


1
Expert's answer
2021-07-06T03:38:09-0400

Given Function :

"F: R\to R , F(x)=x^2"

"i)" The function is not one-one function.

A one-one function is a function in which every element of co-domain has a distinct image in domain.

Here domain as well as co-domain is "R"

"F(x)=y=x^2" .

Here , "y=(-x)^2=x^2"

"y=x^2\implies x=\pm \sqrt{y}"

which means an element in "y=F(x)" has two different elements in domain , "-x" and "x."

Example : "F(x)=y=4" for "x=\pm 2" .

"ii)"

The function is not onto function.

A onto function is a function in which every element of co-domain has at least one image in domain.

Or, Range and Co-domain must be same.

Here domain as well as co-domain is "R" .

But for any negative real number in co-domain will not have image in domain.

"y=x^2\implies x=\pm \sqrt{y}" and "x" is not defined if "y<0" .

So, every element of co-domain has not an image in domain.

Example: "y=-4" "=x^2" has no element in domain.